If x=3+2√2 find I) 1/x, ii) x-1/x, iii)(x-1/x)^3, iv) x^3-2/x^3

Grade:9

1 Answers

Arun
5 years ago
Dear student
 
If x = 3 + 2 sqrt(2)
Now
1/x = 1/ (3 + 2 (sqrt 2)) = 3 - 2 sqrt (2) / (9 -8) = 3 - 2 sqrt(2)
Now
x - 1/x = 4* sqrt (2)
Now
(x-1/x)³ = 64 * 2 * sqrt (2) = 128 * sqrt(2)